### v1

@@ 4 DISCUSSION
This systematic review revealed a growing body of research on AI-assisted development that intersects three key themes: Design, Impact, and Quality. From the Design perspective, the most important areas of research are related to transparency and simplicity of the developing AI tools. Design research focuses primarily on autocompletion and conversational agents, with growing interest in hybrid models. From the Impact perspective, AI improves productivity, especially for repetitive tasks, but also increases verification efforts and risks automation bias. From the Quality perspective, AI-generated code is found to be efficient but prone to errors, security risks, and maintainability issues.
The reviewed studies used qualitative, quantitative, and mixed methods designs, with a considerable part of the research falling outside the qualitative-quantitative division (e.g.,, development of the new AI tools). It should also be noted that 82% of the studies focus on using AI tools in professional settings, while 18% explore AI in educational contexts.
Many reviewed studies focus on how AI affects coding tasks (e.g.,, autocompletion for boilerplate code, conversational agents for complex problem solving) rather than why these tools succeed or fail in particular contexts. Aligning the two perspectives can clarify both the Design of the IDE integrations and the Impact on the practitioners’ everyday workflows (Weber et al., 2024; Robe and Kuttal, 2022). Rather than treating AI as a mere feature, researchers can adopt HCI-informed frameworks that consider user trust, mental models, and organizational culture as central to tool adoption.
It is evident that AI-assisted tools can boost code production speed (Peng et al., 2023; Pandey et al., 2024), but the depth of these gains appears to be strongly context-dependent. Many studies document the promise of conversational agents and autocompletion (Vaithilingam et al., 2023; Ross et al., 2023a), but the success of these paradigms depends on matching the right interaction mode with the right development stage.
The HCI theories on mixed-initiative design (Amershi et al., 2014) recommend interfaces that dynamically adjust the level of AI control, factoring in the user’s expertise and the complexity of the tasks. Novices, for example, may improve their immediate performance yet fail to internalize key programming concepts (Zhou and Li, 2023). Studies of educational contexts (Puryear and Sprint, 2022) suggest that real gains occur when learners receive structured prompting, consistent feedback, and explicit guidance in validating AI outputs (Prather et al., 2024; Kruse et al., 2024). Moreover, advanced developers, while benefiting from speed-ups, must spend a substantial portion of time auditing and refining AI outputs to ensure correctness in highly domain-specific tasks (Mozannar et al., 2024a).
Prior research highlights the risk of subtle flaws, security vulnerabilities, and poor maintainability (Sandoval et al., 2023; Pearce et al., 2021). Despite the existence of automated checks and static analysis, most quality control remains a human responsibility (Mozannar et al., 2024a; Tang et al., 2024). Therefore, inline suggestions, while streamlining coding, warrant the need for transparent AI behavior and meaningful explanations to mitigate “automation bias” (Al Madi, 2022) and acceptance of incorrect code. Effective trust calibration can involve real-time feedback loops (e.g.,, automated test suites) or explicit disclaimers about the AI ’s uncertainty (Kocielnik et al., 2019). In line with HCI work on explainable AI, embedding contextual clues (e.g.,, highlighting AI confidence levels or providing direct links to relevant training data) may calibrate trust more effectively (Kocielnik et al., 2019). Such design elements should also encourage user control, letting developers accept or reject suggestions with a clear sense of the trade-offs (Prather et al., 2024).
The ethical dimensions of AI in software development remain underexamined. For instance, non-native English speakers or beginners may struggle with complex prompts, inadvertently reinforcing existing inequities in open-source communities or corporate teams. Adopting AI within professional software teams shifts norms around collaboration, with some studies likening AI tools to “pair programmers” (Robe and Kuttal, 2022). This conceptual shift raises questions about division of labor and collective ownership of code: developers can offload routine checks to the AI, but may lose deeper engagement with design patterns or system architecture. It also introduces new roles, such as “prompt engineer” or “model tutor”, highlighting the need for training team members not only to code, but also to manage and interpret model outputs (Feng et al., 2024; Cheng et al., 2024b). Integrating AI tools at scale requires policies that balance innovation with accountability. In fields like healthcare or finance, AI-driven coding missteps could carry elevated risk (Pearce et al., 2021). Drawing on sociotechnical systems perspectives in HCI, organizations may need to formalize acceptance criteria for AI-generated suggestions, define clear roles for oversight, and establish audit trails for code provenance (Amershi et al., 2019).
Our findings align with a broader meta-analysis (Vaccaro et al., 2024) of 106 studies on human–AI collaboration, which reveals nuanced performance trade-offs depending on a task type, human expertise, and the relative strengths of humans and AI. The meta-analysis found that human–AI combinations often perform worse than the best individual performer (Hedges’ $g=-0.23$), with performance losses in decision-making tasks and gains in creative content creation. Although these results do not yet seem to be the same for coding tasks in real-world applications, both analyses emphasize the role of user expertise in mediating outcomes. Novices struggle with interpreting AI outputs, often leading to reliance on incorrect suggestions, while experienced developers leverage AI more effectively for repetitive or well-defined tasks. Importantly, both highlight the heterogeneity of impacts, with collaboration being highly task-specific and sensitive to how AI tools align with human workflows. To improve human–AI systems, the meta-analysis suggests balancing human and AI capabilities, while our review emphasizes domain-specific considerations: enhancing explainability, tailoring AI tools to user expertise, and refining prompting mechanisms to reduce errors. These findings jointly point out the importance of designing adaptive, user-centered systems that align with task contexts and user needs to maximize the complementary strengths of humans and AI.
Our systematic review of the literature indicates one recurring issue in the experimental part of the research employed within the HAX field. We observed that studies often feature small sample sizes (median being 17 participants) without justification for the number of participants, and there is a total lack of preregistered research. Given the rapid pace of progress in the area, it is also challenging to assess the potential replicability of ongoing research and to address problems such as publication bias. The replicability crisis has become the most discussed regarding behavioral and social sciences. However, concerns regarding research rigor are also prevalent within Computer Science, particularly in areas such as AI and Machine Learning (Henderson et al., 2019). To address these challenges, we believe that the HAX research community should consider adopting practices from neighboring fields. Particularly psychology, which has faced a replicability crisis but has made significant strides toward overcoming it (Nosek et al., 2022). We propose the widespread adoption of practices such as preregistration, the routine use of statistical power analysis, and the implementation of blinding and randomization procedures where possible. Additionally, embedding transparency checklists as standard practice can guide researchers in documenting their methodology, statistical tests, data exclusion criteria, and data-sharing protocols. By integrating these practices into the peer-review process and rewarding transparency through incentives, the HAX community can align with other fields that have successfully improved replicability and credibility.
@@ 4.1 Threats to the Validity
Sampling Bias: Despite rigorous efforts to include well-known libraries and refine the search string, the possibility of sampling bias remains. Although our approach aimed to minimize bias, the inherent challenge of capturing every relevant work persists. That is why we provide our search protocol to mitigate this threat.
@@
Challenges of industry-related research: It is important to acknowledge that, overall, HAX research faces unique challenges due to its close ties with the industry. One of the most significant challenges of conducting empirical studies in industrial settings is the ”time factor.” Industrial sponsors and participants often have short time horizons. They expect valuable feedback within 1-2 months of research, so that they could make timely business decisions. Although industrial sponsors and project managers are interested in rigorous results, practical constraints sometimes make it challenging to prioritize statistical significance. Analyzing immediate trends and patterns can be sufficient for them to assess the performance of their projects, teams, or technologies. By fostering a collaborative environment, we can develop strategies that balance the demands of industrial collaboration with the standards of academic research.
@@ 5 CONCLUSION
Overall, our findings show that AI-assisted development has already begun transforming the nature of software engineering work, but the journey is far from complete. The core challenges revolve around (1) aligning AI capabilities with human expertise, (2) embedding transparency and accountability into developer tools, and (3) cultivating robust validation practices. From an HCI point of view, these challenges emphasize the importance of user-centered design, inclusive experimentation, and empirically-grounded evaluations that extend beyond short-term productivity metrics. We believe that sustained collaboration among researchers, practitioners, and end users will be essential for guiding innovation, refining best practices, and ultimately unlocking the full potential of AI-assisted development.
@@ Acknowledgements.
The authors have no conflict of interest with the content of this article.
Data availability
All data are available in our supplementary materials Sergeyuk et al. (2025).

### v2

@@ 4 Discussion
We organize the discussion of our systematic literature review of 90 in-IDE HAX studies around topics coverage and gaps (RQ1), implications for practice and research (RQ2), and a future work agenda (RQ3).
RQ1: Extensively studied and under-explored aspects of in-IDE HAX.
The corpus of reviewed works separates cleanly into two contexts of use for AI in software development, namely professional (71/90) and educational (19/90). Within professional settings, most papers examine practical use in coding and maintenance, and a large share evaluate assistance without situating it in a specific SDLC stage. Educational papers focus on how learners use assistance to solve programming problems. Here as well, stage information is often omitted, which limits comparison across contexts. This weakens interpretability because the same interface may have different effects at requirements, implementation, testing, or evolution. Future work should report the stage explicitly and use a simple task taxonomy that allows cross-paper synthesis. Task design employed in studies evolves over time. Early studies rely on small algorithmic exercises that resemble simple games and interview-style problems. More recent work introduces project-level analysis of workflows and begins to include tasks related to documentation, testing, debugging, and migration. This broadening improves ecological validity, yet it also increases the number of papers that label the task as unspecified. Clearer task descriptions and stable taxonomies are needed to preserve comparability as the scope widens.
Several papers frame  AI-assistance as a form of pair programming in which the system acts as a collaborator rather than a tool. This view aligns with the observed move from isolated completions to co-creative workflows. It also raises concrete needs in both contexts, namely training in model interaction and oversight in professional teams, and scaffolds that keep learners in control in educational settings. In summary, the evidence is richest for professional coding and maintenance, with a growing variety in task types and a notable reporting gap on the SDLC stage. The next step is to include non-adopters, expand coverage to early and late stages, and standardize task and stage reporting so that results can be compared across settings and over time.
RQ2: Key findings and implications in the field of HAX in IDEs.
Impact. Across reviewed studies, there is evidence that AI assistance in IDEs changes how developers create and verify code. The main benefit is faster progress on routine or well-scoped work, often explained by fewer context switches and reduced need for boilerplate code writing. The main limit appears on tasks that embed proprietary knowledge or complex logic, where current systems struggle to capture project-specific nuance (Peng et al., 2023; Pandey et al., 2024; Khemka and Houck, 2024). Code verification is, therefore, central. About fifty percent of developers’ time is now spent on inspecting suggestions, refining prompts, and repairing code (Mozannar et al., 2024a). These findings argue for budgeting verification explicitly and for evaluating impact with edits and verification effort in addition to the suggestions’ acceptance rate and time spent on task. Effects are heterogeneous across populations. Novices complete tasks more quickly and report lower workload when the use of in-IDE AI is structured, yet they are vulnerable to subtle errors and may overtrust suggestions (Kazemitabaar et al., 2023a; Prather et al., 2024; Zhou and Li, 2023). Professionals benefit from speed on familiar patterns but may under-rely on AI and reject correct suggestions if it lacks a clear rationale (Sun et al., 2022). Trust also varies by context. It decreases in high-stakes and testing contexts and increases for routine or proof-of-concept work, which argues for visible boundaries and rationale at decision time (Wang et al., 2024; Brown et al., 2024).
Design. Reviewed studies converge on three core design principles for AI in IDEs, namely rich context awareness, explainability with transparency, and user control with adaptability. These principles guide how AI systems should be built. Moreover, it is highlighted that the two dominant interaction modes, autocompletion and conversational, should work together. A smooth switch between them, a shared context pool, and consistent state across views allow developers to move from local edits to higher-level reasoning without friction (Weber et al., 2024; Yen et al., 2023). As conversational interaction with AI enters IDEs, prompting becomes an important learnable activity. Structured prompts and reusable prompt-outcome pairs align generation with intent and support calibrated acceptance (Kruse et al., 2024; Denny et al., 2023; Fagadau et al., 2024). Interfaces that ask the user to state the next step before code is shown and that supply brief rationales further stabilize decisions (Kazemitabaar et al., 2025), as well as step-by-step explanations of the generated code for code validation and refinement Tian et al. (2023, 2024). Explanation and control must be available at decision time. Useful aids in verification include surfacing runtime values, side-by-side alternatives, retrieval of reference examples, static analysis, and unit tests inside the editor Ferdowsi et al. (2024); Pearce et al. (2022); Weisz et al. (2022); Vaithilingam et al. (2022). Context-aware explanations and explicit scope controls make it clear where and why the assistant acts (Vaithilingam et al., 2023; Wang et al., 2024; Yan et al., 2024). When available, calibrated uncertainty cues or predicted edit likelihood should direct review effort, without relying on raw generation probabilities alone (Spiess et al., 2025; Vasconcelos et al., 2025). In educational settings, structured prompts, brief rationales, staged contribution, and quick checks inside the IDE help learners stay in control (Robe and Kuttal, 2022; Kazemitabaar et al., 2024b, 2023a; Prather et al., 2024; Zhou and Li, 2023). Embodied or mixed reality agents are useful when the goal is to practice collaboration routines with feedback tied to the editor context (Manfredi et al., 2023).
Quality. Studies in this direction examine correctness, security, readability, and maintainability of code produced with AI assistance inside the IDE. Across these papers, the central pattern is a trade-off between speed and assurance. Assistance can accelerate progress, yet LLM-generated code that passes superficial checks may still contain logical flaws or insecure patterns, which shifts effort from authoring toward checking (Wang et al., 2025; Nguyen and Nadi, 2022; Yetistiren et al., 2022; Al Madi, 2022; Asare et al., 2023; Sandoval et al., 2023; Asare et al., 2024; Mozannar et al., 2024a; Izadi et al., 2024). LLM as judge signals can complement human review for correctness (Tong and Zhang, 2024), but correctness outcomes depend not only on model properties but also on developer practice, including how prompts are written, how suggestions are inspected, and how teams organize review (Kruse et al., 2024; Feng et al., 2024; Kuang et al., 2024; Mozannar et al., 2024a). Readability and maintainability concerns recur alongside correctness. Short and compact suggestions reduce typing but can hinder clarity of intent through unconventional names, missing comments, or terse structure, which complicates later edits and team review (Nguyen and Nadi, 2022; Yetistiren et al., 2022; Vaithilingam et al., 2022; Al Madi, 2022). Explanatory aids and refactoring support inside the editor help address these issues and make suggested changes easier to justify (Yan et al., 2024; Ferdowsi et al., 2024). Security is a shared responsibility as well. Models can reproduce unsafe templates from training data, while developers may accept risky code under time pressure (Pearce et al., 2022; Sandoval et al., 2023; Asare et al., 2023, 2024; Al Madi, 2022; Mozannar et al., 2024a; Vaithilingam et al., 2022; Gardella et al., 2024). Taken together, these findings call again for verification with tests and static analysis inside the editor, retrieval of relevant examples, and concise explanations at decision time, so that speed and assurance can coexist (Vaithilingam et al., 2022; Liu et al., 2024; Yan et al., 2024; Ferdowsi et al., 2024).
RQ3: Future research and development agenda.
Analysis of the future work statements yields a ranked set of topics and methodologies to study in-IDE HAX. The most frequent cluster concerns productivity factors $(43)$. Here, the field should move from isolated gains to comparative and larger-scale evaluations that report acceptance, edits, and verification effort in addition to time. The next clusters focus on the design of assistance $(29)$ and audits of AI generated code $(28)$. For design, work should advance hybrid workflows that integrate autocompletion and conversation, with prompt scaffolds and explanation layers tied to decision points (Yen et al., 2023; Ross et al., 2023a; Brown et al., 2024; Guo et al., 2024). For audit, the field needs targeted test suites, stronger security coverage, and longitudinal assets that track change over time (Agarwal et al., 2024; Mozannar et al., 2024c; Asare et al., 2024; Pearce et al., 2022; Asare et al., 2023; Ságodi et al., 2024). Prompting support appears in nineteen statements and should prioritize guided decomposition and hierarchical specification, with interactive explanations that support validation and refinement (Yen et al., 2023; Brown et al., 2024). Skill building with AI $(17)$ and education with AI $(16)$ call for structured practice that protects learning goals inside authentic IDE settings (Prather et al., 2023; Zhou and Li, 2023; Angert et al., 2023; Gardella et al., 2024; Ferdowsi et al., 2024). Personalization is equally frequent $(16)$ and should aim at policies that adapt timing, modality, scope, and context to user profile and task state, with simple controls for frequency and snooze (Vaithilingam et al., 2023; Ross et al., 2023a; de Moor et al., 2024; Chopra et al., 2024; Koohestani and Izadi, 2025). Explainability and transparency are requested by fifteen papers and should be evaluated for their effect on acceptance. Verification support $(13)$ and broader SDLC coverage $(12)$ align with the observed gaps in stage specification and with the shift of time toward checking. Less frequent but still important topics for study are trust and reliability $(11)$, IDE redesign for AI $(10)$, assistant context enrichment $(6)$, user mental models, proactivity $(4)$, user control $(3)$, and governance $(3)$. These numbers do not imply that lower frequency topics are unimportant. They indicate where evidence and assets are thin and where careful study can yield high value.
Methodological and open science recommendations.
Across the corpus, sample sizes are often underpowered and rarely justified. The median is seventeen participants. According to prior studies, for HCI practice, the most common sample size is twelve and the median is eighteen; in-person studies have a median of fifteen, while remote studies have a median of seventy-seven (Caine, 2016). Although many quantitative studies at these sizes will be underpowered and therefore benefit from replication. We recommend a priori power analysis for confirmatory comparisons and saturation arguments for qualitative designs, with explicit reporting of recruitment constraints. Collaboration between academia and industry can improve recruitment and enable longitudinal assets. To support replication, we recommend sharing versioned repositories with study materials and data where feasible, including codebooks, prompts, task descriptions, analysis scripts, and deidentified datasets with clear metadata and a data dictionary. We also recommend using preregistration, blinding, and randomization when appropriate, and embedding transparency checklists that record methodology, exclusion criteria, and data sharing plans.
Takeaways.
Treat verification as a main part of the interaction with AI assistants, keeping tests, static analysis, and reference examples inside the editor and by surfacing runtime evidence to support calibrated acceptance decisions. Evaluate the effect beyond time spent on the task by reporting acceptance, edits, and verification effort. Design AI assistance as a hybrid of autocompletion and conversation that shares context, offers brief explanations, and scope controls at decision time. Structure educational use of AI in IDEs with staged contribution of AI, stepwise reveal of its suggestions, and controls for their quick checks. For studies, report the SDLC stage and expand coverage of tasks completed with AI to early and late stages. Strengthen methodological rigor through justified sample sizes, preregistration where appropriate, transparent reporting, and collaboration with industry for larger, ecologically valid, and longitudinal evaluations.
@@ 4.1 Threats to the Validity
Sampling Bias: Despite efforts to include well-known libraries and refine the search string, the possibility of sampling bias remains. Our recall-oriented query design can bias the pool toward professional contexts, since educational deployments do not always occur inside full IDEs or are indexed under alternative platforms. We mitigated this by including multiple synonyms where supported and by manually screening every record for the inclusion criterion. Although our approach aimed to minimize bias, the inherent challenge of capturing every relevant work persists. That is why we provide our search protocol to mitigate this threat.
@@
Challenges of industry-related research: It is important to acknowledge that, overall, in-IDE HAX research faces unique challenges due to its close ties with the industry and professional context. One of the most significant challenges of conducting empirical studies in industrial settings is the "time factor". Industrial sponsors and participants often have short time horizons. They expect valuable feedback within 1-2 months of research, so that they could make timely business decisions. Although industrial sponsors and project managers are interested in rigorous results, practical constraints sometimes make it challenging to prioritize statistical significance. Analyzing immediate trends and patterns can be sufficient for them to assess the performance of their projects, teams, or technologies. By fostering a collaborative environment, we can develop strategies that balance the demands of industrial collaboration with the standards of academic research.
@@ 5 CONCLUSION
This review offers a structured synthesis of current research on AI-powered assistance in integrated development environments. Drawing on 90 studies, we identified dominant trends, conceptual gaps, and methodological characteristics of the emerging field of in-IDE HAX.
We found that research to date focuses heavily on a small set of tools, most often GitHub Copilot, and concentrates primarily on the implementation stage of the SDLC. While this has produced early insights into how developers engage with AI assistance during code writing and modification, it leaves other critical stages such as requirements analysis, testing, and deployment underexplored. To make progress, the field must broaden its lens. Future research should investigate diverse AI assistants across varied development environments, extend coverage to earlier and later lifecycle stages, and prioritize longitudinal and comparative designs that capture how practices and their impact evolve over time. Attention is also needed to support personalization and trust calibration, reduce validation burden, and address broader quality and governance concerns. Finally, we note that methodological rigor remains an area for improvement. Methodologically, studies tend to prioritize short-term evaluations, often with underpowered samples and limited contextual variation. As the field matures, advancing cumulative evidence will require the adoption of robust empirical practices and closer collaboration across research groups.
In sum, the reviewed literature demonstrates growing interest in the integration of AI into developer tools but also highlights the need for deeper empirical grounding, broader tool coverage, and stronger methodological foundations. Addressing these challenges will be essential for supporting effective, transparent, and user-centered AI assistance in software development workflows.
@@ Acknowledgements.
Declarations
Funding
This research was supported by JetBrains Research.
Ethical approval
Not applicable.
Informed consent
Not applicable.
Author Contributions
Agnia Sergeyuk, Ilya Zakharov, Ekaterina Koshchenko: Conceptualization, methodology, literature survey, writing, original draft, and editing.
Maliheh Izadi: Conceptualization, methodology, writing, original draft, and review.
Data availability
All data are available in our supplementary materials (Sergeyuk et al., 2025).
@@
The authors declare that they have no conflict of interest.
Clinical Trial Number
Not applicable.